# A novel numerical framework for three-dimensional fully resolved simulation of freely falling particles of arbitrary shape

### v1

@@ 3 Results
The numerical free-fall experiment is validated in the following sections at different Reynolds numbers for particles of spherical, spheroidal, and a lapilli-sized volcanic particles. The numerical parameters of the simulations are summarized in Table 2. $N$ is defined as the number of lattice cells used to resolve a particle diameter $d$, where $d$ the longest axis for the ellipsoids. In the case of the arbitrary-shaped volcanic particle, $d$ is the diameter of an equivalent circle of area equal to the cross-sectional are of the broadest particle side. The discrete cell spacing is defined as $\delta x=d/N$. The resolution The time discretization $\delta t$ is fixed by fixing the terminal velocity (analytical, empirical, or experimental prediction) to a value of 0.02 in lattice units.
Physical parameters Numerical parameters Results $d$ [mm] $\kappa$ $\nu$ [$\mathrm{m^{2}/s}$] $el$ $fl$ $w^{*}$ $h^{*}$ $h_{\mathrm{ini}}^{*}$ $h_{\mathrm{end}}^{*}$ $l_{\mathrm{sponge}}^{*}$ $N$ $v_{\mathrm{T}}$ [$\mathrm{m/s}$] $\mathrm{Re}_{T}$ Sphere-1 0.148 996.26 1.51e-5 1.0 1.0 10.0 40.0 35.0 5.0 4.0 400 0.5233 5.12 Sphere-2 0.140 996.26 1.51e-5 1.0 1.0 10.0 40.0 35.0 5.0 4.0 300 0.4848 4.49 Spheroid-1 0.140 996.26 1.51e-5 1.0 0.2 17.1 51.3 46.2 5.1 4.1 550 0.3405 3.15 Spheroid-2 0.140 996.26 1.51e-5 0.2 1.0 29.3 87.9 79.1 8.8 7.0 550 0.3630 3.36
Table 2: Parametrization and summary of results of the virtual free-fall numerical experiments. For simulations with multiple resolution for the same particle, the value of $N$, $v_{\mathrm{T}}$, and $\mathrm{Re}_{T}$ correspond to the simulation with the highest resolution in this table. The right column shows the average terminal velocity $v_{\mathrm{T}}$ reached in the statistically stationary state, as well as the associated Reynolds number $\mathrm{Re}_{T}$.
3.1 Grid convergence
A grid convergence study was conducted at a non-vanishing Reynolds number close to $\mathrm{Re}=5$, using the numerical parameter series ‘Sphere-1’ in Table 2. The study demonstrates a convergence of the drag force on a sphere as a function of grid resolution, and a comparison with the empirical drag predicted in [67] shows a satisfactory convergence of the total force already at a low resolution of 10 grid points (Figure 2). A decomposition into force contributions however shows that pressure force and the shear force, taken individually, are inaccurate at low resolution and start showing signs of grid convergence at a higher resolution of 30 grid points only.
Figure 2: Convergence of the total drag and the relative contributions of pressure and viscous forces as a function of grid resolution.
@@ 3.1 Grid convergence
This study shows that sufficient grid convergence is achieved at values of resolution starting at $N=30$. It must be emphasized that proper numerical values of the the pressure and shear components of the force are critically important to model the free-fall behavior, and especially the onset of rotational motion, of particles. While it can be tempting to choose a grid resolution on ground of a convergence study of the drag force alone, it is now clear that this would lead to under-resolved simulations. In conclusion, a resolution of $N=30$ for most of the simulations of finite, low Reynolds number presented below.
3.2 Rotating sphere at low Reynolds number
@@
For the validation of sphere motion in inertial flow at different Reynolds numbers, we compare it to the empirical relationship proposed by Clift & Gauvin [67]. This model presents the standard drag curve throughout the transitional and Newton regimes ($\mathrm{Re}<10^{5}$):
@@
Through time-integration of Newton’s equations of motion, this drag curve translates into a free-fall trajectory of a spherical particle from rest to terminal velocity $v_{\mathrm{T}}$. The equation is solved through an implicit Euler scheme, where the time evolution of the sphere velocity is given by the following Basset-Boussinesque-Oseen equation [68]:
A solution for rotational motion was first considered by Basset [69] and later improved by Feuillebois and Lasek [70]. Here, a constant torque $T$ is applied to a sphere initially at rest. The sphere begins to accelerate angularly and ultimately reaches the terminal angular velocity $\omega_{\mathrm{st}}$, given by
$\omega_{\mathrm{st}}=\frac{T}{m_{\mathrm{sphere}}}\frac{\kappa}{6\nu},$ (33)
where $m_{\mathrm{sphere}}$ is the mass of the sphere and $\kappa$ is the density ratio between the sphere and the fluid. The time evolution of the angular velocity $\omega$ is described by the relation
For a sphere with homogeneous mass distribution, the moment of inertia $I$ takes the value $I=2/5m_{\mathrm{sphere}}r^{2}$, and the term $\rho_{\mathrm{f}}/I$ in the equation above is given by
$\frac{\rho_{\mathrm{f}}}{I}=\frac{15}{8}\frac{1}{\kappa r^{5}\pi}.$ (35)
@@
For the DNS, we considered the equivalent problem of a sphere which, in absence of an external torque, starts rotation under the effect of an ambient linear shear flow. The far-field flow velocity is defined as
$v_{x}=0,\quad v_{y}=v_{0}+Gy,\quad v_{z}=0,$ (36)
@@
where the coordinates $x$, $y$, and $z$ represent a fixed reference frame. Here, $G$ is the shear rate of the ambient flow, and we denote the rotation rate of the ambient flow as $\omega_{f}=G/2$. As pointed out for example in [71], the terminal angular velocity of a sphere is equal to the ambient flow rate in a Stokes flow, $\omega_{f}=\omega_{\mathrm{st}}$, while a slip angular velocity is observed at finite Reynolds numbers. We define the Reynolds number in a rotational flow as
$\mathrm{Re}_{G}=\frac{4Gr^{2}}{\nu},$ (37)
Leading to the following expression for the torque term in Eq. 34:
$\frac{T}{\rho_{\mathrm{f}}}=\mathrm{Re}_{G}\frac{9}{16}r\pi\nu^{2}.$ (38)
For the DNS, the Reynolds number is fixed to a finite value of $\mathrm{Re}\sim 5$, with simulation parameters as $d=0.140$ mm, $\kappa=1000$, $\nu=1.73e^{-5}\mathrm{m^{2}/s}$, $w^{*}=20$, $h^{*}=30$, $h_{\mathrm{ini}}^{*}=20$, $h_{\mathrm{end}}^{*}=10$, $l_{\mathrm{sponge}}^{*}=3$, and $N=300$. For the rotational motion, a constant torque is applied to the sphere, gravity is disabled, and a no-slip condition without sponge zone is applied on all walls. The numerical results for translational and rotational motion are shown in Figure 3, and compared to the empirical model relations 32 and 34. The simulations show a good match with the analytical curves, although for the translational motion, the terminal velocity is reached slightly faster in the DNS than in the analytical curve. This observation is compatible with the results of a spectral method obtained at a density ratio $\kappa=1.05$ in [72]. As explained further in [72], the discrepancy between analytical and numerical solution is mainly due to the effect of finite Reynolds number. For the rotational motion, the numerical solution matches the analytical one closely, with an cumulative error that builds up due to a long time range spanning over two orders of magnitude.
Figure 3: Free translation (left curve) and free rotation (right curve) in a Stokes flow.
3.3 Terminal velocity of sphere at different Reynolds number
As a next validation case, we measured the terminal velocity $v_{\mathrm{T}}$ of a freely falling sphere. This test is more challenging than a simple numerical evaluation of the sphere drag, as the spheres are moving and rotating and traverse the cells of the fluid grid. The DNS domain boundaries had a fixed extent of $w^{*}=10$ and $h^{*}=30$, and the same resolution of $N=300$. However, depending on the particle Reynolds number $\mathrm{Re}$, $h_{\mathrm{ini}}^{*}=27.0$, $h_{\mathrm{end}}^{*}=3.0$ and $l_{\mathrm{sponge}}^{*}=2.4$ are used when $\mathrm{Re}<30$; $h_{\mathrm{ini}}^{*}=23.0$, $h_{\mathrm{end}}^{*}=3.0$ and $l_{\mathrm{sponge}}^{*}=2.4$ for $100020\leq\mathrm{Re}<$ 1000; and $h_{\mathrm{ini}}^{*}=20.0$, $h_{\mathrm{end}}^{*}=3.0$ and $l_{\mathrm{sponge}}^{*}=2.4$ for $\mathrm{Re}\geq 1000$. The results are summarized in Figure 4 and in Table 3. It can be seen that in a laminar regime, the numerical terminal velocity matches the empirical value with a two to three-digit accuracy, and for higher Reynolds number up to $50,000$, the simulation results remain within a $6\%$ difference from the empirical one.
Figure 4: Free-fall velocity of spheres at various $\mathrm{Re}$, compared to empirical model prediction in Eq. 31.
$\mathrm{Re}^{\mathrm{ref}}$ 0.57* 1.36 4.57 5.24 6.34 29.34 129.40 586.36
$d$ [mm] 0.062 0.088 0.140 0.148 0.160 0.315 0.659 1.488
$v_{\mathrm{T}}^{\mathrm{ref}}$ [m/s] 0.138* 0.234 0.493 0.535 0.599 1.408 2.968 5.956
$v_{\mathrm{T}}^{D}$ [m/s] 0.142 0.239 0.485 0.523 0.584 1.358 2.933 6.035
$\Delta v_{\mathrm{T}}$ [%] 2.82 2.09 -1.65 -2.29 -2.57 -3.68 -1.19 1.31
$p^{*}$ [%] 35.80 38.76 42.67 43.10 43.46 48.37 57.50 68.90
@@ 3.3 Terminal velocity of sphere at different Reynolds number
Table 3: Evolution of terminal velocities of spheres at different Reynolds number. For reference, a empirical model of [67] is used to approximate the drag force in Eq. 31, and the resulting terminal velocity $v_{\mathrm{T}}^{\mathrm{ref}}$, and the Reynolds number $\mathrm{Re}^{\mathrm{ref}}$. For the low $\mathrm{Re}^{\mathrm{ref}}$ case, the reference terminal velocity is computed from Stokes’ law: $v_{\mathrm{T}}=(1/18)(\kappa gd^{2}/\nu)$. The terminal velocity obtained from the DNS is shown as $v_{\mathrm{T}}^{D}$, $\Delta v_{\mathrm{T}}=(v_{\mathrm{T}}^{D}-v_{\mathrm{T}}^{\mathrm{ref}})/v_{\mathrm{T}}^{D}$, and $p^{*}$ indicate the contribution of the pressure force to the total drag force at the steady-state. The mass-densities of the sphere and the fluid are $\rho_{\mathrm{p}}=1200$ and $\rho_{\mathrm{f}}=1.2045$ kg/m3 respectively, and the kinematic viscosity of the fluid is $\nu=1.51\times 10^{-5}$ m2/s, representative of the air at 20oC at the mean sea level.
For validating the time history of settling velocity from the numerical free-fall simulation, a Reynolds of $Re=4.49$ is chosen, with parameters summarized under the label ‘Sphere-2’ in Table 2. Figure 5 compares the obtained numerical and empirical curve of Eq. 31.
Figure 5: Time-evolution of sphere free-fall at finite Reynolds number $Re=4.5$. Comparison between numerical and empirical values.
A very close match is obtained for the particle time-evolution, with an expected discrepancy of $2\%-3\%$ regarding the terminal velocity.
3.4 Spheroids

### v2

@@ 3 Results
The numerical free-fall experiment is validated in the following subsections for spherical, spheroidal, and an irregular volcanic particles at different Reynolds numbers. The numerical parameters of the Palabos Turret simulations are summarized in Table 2 or in corresponding results subsections for different cases considered in this study, where the parameters with * are dimensionless. The discrete cell spacing is defined as $\delta x=w/N$. The time step $\delta t$ is defined by normalizing the terminal velocity (analytical, empirical, or experimental prediction) to a value of 0.02 in lattice units. In all simulations, gravity $g$ has a value of 9.81 m/s2. The Palabos Turret simulations are performed in the parallelized super-computing facilities of the Max Planck Society. The results of the Palabos Turret simulations are compared with analytical models and experimental measurements performed with the Göttingen Turret [16] and the UNIGE Vertical Wind Tunnel [64, 10]. The settling and terminal velocities presented in the results are the absolute of the particle velocity component in the $y$ direction, i.e. $|v_{y}|$. The terminal velocity $v_{\mathrm{T}}$ is defined as $|v_{y}|$ at the point in time at which $|a_{y}|$ falls below 0.01 m/s2.
3.1 Domain and mesh independence tests
[…]
@@ 3.1 Domain and mesh independence tests
A subsequent grid convergence study was conducted at a non-vanishing Reynolds number $\mathrm{Re}\sim 5$, using the numerical parameters of ‘Sphere-1’ in Table 2. The study demonstrates a convergence of the drag force $F_{D}$ on a sphere to its gravitational force $F_{g}$ as a function of the grid resolution (Figure 2(a)), and the relative difference between $F_{g}$ and the terminal state $F_{D}$ is already 0.7% even at a low resolution of $N=100$ or $\delta x=d/10$. A decomposition into force contributions, however, shows that the pressure force and the shear / viscous force, when considered individually, are inaccurate at low resolution and start showing signs of grid convergence at a higher resolution of $N=300$ only. It must be emphasized that proper numerical values of the the pressure and shear components of the force are critically important to model the free-fall behavior, and especially for the onset of rotational motion of the particles. While it can be tempting to choose a grid resolution on ground of a convergence study of the drag force alone, it is clear that this would lead to under-resolved simulations.
Figure 2: (a) Impact of grid resolution $N$ on the convergence of drag force $F_{D}$ to the gravitational force $F_{g}=mg$, and its pressure and shear / viscous force contributions. Here $m$ is the mass of the particle. The values are the percentage of the pressure force to the total drag force. At the Stokes regime $\mathrm{Re}<<1$, the hydrodynamic pressure forces on the particle surface contributes 33% to the terminal drag force [65]. (b) Validation of the time series of settling velocity of a sphere during free-fall (left) and free-rotation (right) with the empirical Eqs. 36 and 37. Time $t$ and settling velocity $v$ is non-dimensionalized as $t^{*}=t*v_{\mathrm{T}}/d$ and $v^{*}=v/v_{\mathrm{T}}$. $v_{\mathrm{T}}$ is the terminal velocity from the Palabos Turret simulation. The curves from the free-rotation Palabos Turret simulation is shifted in time for better visibility.
Physical parameters Numerical parameters Results $d$ [mm] $\kappa$ $\nu$ [$\mathrm{m^{2}/s}$] $el$ $fl$ $w^{*}$ $h^{*}$ $h_{\mathrm{ini}}^{*}$ $h_{\mathrm{end}}^{*}$ $l_{\mathrm{sponge}}^{*}$ $N$ $v_{\mathrm{T}}$ [$\mathrm{m/s}$] $\mathrm{Re}$ Sphere-1 0.148 996.26 1.51e-5 1.0 1.0 10.0 40.0 35.0 5.0 4.0 400 0.5233 5.12 Sphere-2 0.140 996.26 1.51e-5 1.0 1.0 10.0 40.0 35.0 5.0 4.0 300 0.4848 4.49 Spheroid-1 0.140 996.26 1.51e-5 1.0 0.2 17.1 51.3 46.2 5.1 4.1 550 0.3405 3.15 Spheroid-2 0.140 996.26 1.51e-5 0.2 1.0 29.3 87.9 79.1 8.8 7.0 550 0.3630 3.36
Table 2: Simulation parameters and a summary of results of the virtual free-fall Palabos Turret simulations. $d$ is the diameter of a volume-equivalent sphere, $\kappa=\rho_{\mathrm{p}}/\rho_{\mathrm{f}}$ is the particle to fluid density ratio, $\nu$ is the kinematic viscosity of fluid, and particle shape parameters, such as elongation $el=b/a$ and flatness $fl=c/b$ are defined as the ratios of the largest $a$, intermediate $b$ and smallest $c$ dimensions of the particle, orthogonal to each other. Numerical length parameters are non-dimensionalized with respect to $d$, e.g. $w^{*}=w/d$, $h^{*}=h/d$. For simulations with multiple resolution for the same particle, the value of resolution $N$, terminal velocity $v_{\mathrm{T}}$, and particle Reynolds number $\mathrm{Re}$ here correspond to the simulation with the highest resolution. The last two columns show the average $v_{\mathrm{T}}$ in the statistically stationary state, as well as the associated $\mathrm{Re}$.
3.2 Transient motion of a sphere at low Reynolds number
@@
For the validation of sphere motion in inertial flow at different Reynolds numbers, we compare it to the empirical relationship proposed by Clift & Gauvin (1971) [66]. This model presents the standard drag curve throughout the transitional and Newton regimes ($\mathrm{Re}<10^{5}$):
@@
Through time-integration of Newton’s equations of motion, this drag curve translates into a free-fall trajectory of a spherical particle from rest to terminal velocity $v_{\mathrm{T}}$. The equation is solved through an implicit Euler scheme, where the time evolution of the sphere velocity is given by the Basset-Boussinesque-Oseen equation 36 [67]. For validating the time history of settling velocity from the free-fall Palabos Turret simulation, a Reynolds of $Re=4.49$ is chosen, with parameters summarized under the label ‘Sphere-2’ in Table 2. Figure 2(b) compares the obtained numerical result and empirical curve from Eqs. 31 and 36. The time series of the free-fall settling velocity from the Palabos Turret simulation is in excellent agreement with the empirical curve, and shows a relative difference of 1.7% when comparing the terminal velocities.
@@
For the validation of sphere motion in a shear flow with no gravity, we simulated a sphere that starts to rotate without external torque under the effect of a linear shear flow in the ambient. The far-field flow velocity is defined as
$v_{x}=0,\quad v_{y}=v_{0}+Gy,\quad v_{z}=0,$
@@
where the coordinates $x$, $y$, and $z$ represent a fixed reference frame. Here, $G$ is the shear rate of the ambient flow, and we denote the rotation rate of the ambient flow as $\omega_{f}=G/2$. An analytical formulation for the rational motion of a sphere in inertial flow is detailed in B. As pointed out for example in [68], the terminal angular velocity of a sphere is equal to the ambient flow rate in a Stokes flow, $\omega_{f}=\omega_{\mathrm{st}}$, while a slip angular velocity is observed at finite Reynolds numbers. We define the Reynolds number in a rotational flow as $\mathrm{Re}_{G}=4Gr^{2}/\nu$, where $r$ is the radius of the sphere. The torque term in Eq. 37 is therefore $T=9\mathrm{Re}_{G}r\pi\nu^{2}\rho_{\mathrm{f}}/16$. In the Palabos Turret simulation, the Reynolds number is fixed to a finite value of $\mathrm{Re}\sim 5$, with simulation parameters as $d=0.140$ mm, $\kappa=1000$, $\nu=$ 1.73e-5 m2/s, $w^{*}=20$, $h^{*}=30$, $h_{\mathrm{ini}}^{*}=20$, $h_{\mathrm{end}}^{*}=10$, $l_{\mathrm{sponge}}^{*}=3$, and $N=300$. For the rotational motion, the gravity is disabled, and a no-slip boundary condition without sponge zone is applied on all walls. The numerical result of rotational motion is shown in Figure 2(b), and compared to the empirical model Eq. 37. The time series of the free-rotation settling velocity from the Palabos Turret simulation is in good agreement with the empirical curve despite the sphere is simulated at a Reynolds number beyond the Stokes regime. The relative difference in the terminal velocities from the Palabos Turret simulation and the analytical model is 4%.
3.3 Terminal velocities of spheres at different Reynolds number
For the next set of validation, we compared the terminal velocities $v_{\mathrm{T}}$ of freely falling spheres of different sizes, thus different Reynolds numbers. For this case, heavy spheres of density $\rho_{\mathrm{p}}=1200$ kg/m3 are simulated to fall in air of density $\rho_{\mathrm{f}}=1.2045$ kg/m3, resulting to a high particle to fluid density ratio. Such simulations are more challenging than a simple numerical evaluation of the sphere drag. At such high density ratio, these spheres exhibit higher settling velocity and the algorithm needs to accommodate the rotational and translational motion of the solid body in the fluid grid. Moreover, with an increase in sphere Reynolds number $\mathrm{Re}$, the normalized vertical free-fall distance to reach the terminal state increases exponentially. This results in an exponential increase in the computational time to reach the terminal velocity at higher $\mathrm{Re}$.
The Palabos Turret simulation domain boundaries for this validation had a fixed extent of $w^{*}=10$ and $h^{*}=30$. However, depending on the $\mathrm{Re}$, $h_{\mathrm{ini}}^{*}$, $h_{\mathrm{end}}^{*}$ and $l_{\mathrm{sponge}}^{*}$ are varied to accommodate the gradually increasing wake region behind the sphere as the Reynolds number increases. When $\mathrm{Re}<30$, $h_{\mathrm{ini}}^{*}=27.0$, $h_{\mathrm{end}}^{*}=3.0$ and $l_{\mathrm{sponge}}^{*}=2.4$ are used with a grid resolution of $N=300$. When $100020\leq\mathrm{Re}<$ 1000, $h_{\mathrm{ini}}^{*}=23.0$, $h_{\mathrm{end}}^{*}=3.0$ and $l_{\mathrm{sponge}}^{*}=2.4$ are used with the same resolution of $N=300$. And when $\mathrm{Re}\geq 1000$, $h_{\mathrm{ini}}^{*}=20.0$, $h_{\mathrm{end}}^{*}=3.0$ and $l_{\mathrm{sponge}}^{*}=2.4$ are used with an increased resolution of $N=400$ to accommodate for the thinning of the boundary layers near the sphere surface and for capturing the small scale vortical structures in the wake region.
$d$ [mm] 0.062 0.088 0.140 0.148 0.160 0.315 0.659 1.488 3.770 10.530
$v_{\mathrm{T}}^{D}$ [m/s] 0.142 0.239 0.485 0.524 0.584 1.358 2.933 6.035 10.29 17.394
$\mathrm{Re}$ 0.58 1.39 4.49 5.13 6.18 28.30 127.88 594.13 2566.6 12118
$\Delta F$ [%] 0.01 0.24 0.28 0.19 0.48 0.72 0.96 1.04 1.48 3.12
$v_{\mathrm{T}}^{R}$ [m/s] 0.138 0.237 0.488 0.529 0.590 1.382 2.952 5.993 10.891 17.736
$\Delta v_{\mathrm{T}}$ [%] 2.82 0.69 -0.71 -0.91 -1.09 -1.75 -0.65 0.70 -5.84 -1.97
$p^{*}$ [%] 35.80 38.76 42.67 43.10 43.46 48.37 57.50 68.90 88.20 97.09
@@ 3.3 Terminal velocities of spheres at different Reynolds number
Table 3: Evolution of the terminal velocities of freely falling spheres for $5.8\times 10^{-1}<\mathrm{Re}<1.2\times 10^{4}$. The terminal velocity $v_{\mathrm{T}}^{D}$ obtained from the Palabos Turret simulation is used to calculate particle Reynolds number as $\mathrm{Re}=dv_{\mathrm{T}}^{D}/\nu$. The kinematic viscosity of the fluid $\nu$ is $1.51\times 10^{-5}$ m2/s. The relative difference of the gravitational force $F_{g}=mg$ and terminal state drag force $F_{D}$ from the Palabos Turret simulation is shown as $\Delta F=(F_{g}-F_{D})/F_{g}$. Based on $\mathrm{Re}$, a empirical model [66] is used to approximate the drag coefficient $C_{D}$ (Eq. 31), and a reference terminal velocity is obtained as $v_{\mathrm{T}}^{R}=(8F_{g}/(\pi d^{2}\rho_{\mathrm{f}}C_{D}))^{0.5}$. The density of the air $\rho_{\mathrm{f}}$ is $1.2045$ kg/m3, where $\nu$ and $\rho_{\mathrm{f}}$ are representative of the air at 20oC at the mean sea level. For the low $\mathrm{Re}$ case ($\mathrm{Re}=0.58$), the $v_{\mathrm{T}}^{R}$ is computed from Stokes’ law as $\rho_{\mathrm{p}}gd^{2}/(18\rho_{\mathrm{f}}\nu)$. The density of the sphere $\rho_{\mathrm{p}}$ is $1200$ kg/m3. The relative difference of velocities is $\Delta v_{\mathrm{T}}=(v_{\mathrm{T}}^{D}-v_{\mathrm{T}}^{R})/v_{\mathrm{T}}^{D}$. $p^{*}$ indicate the contribution of the pressure force to the total drag force at the terminal state obtained from the stress tensor computed in the Palabos Turret simulations.
The results are summarized in Table 3 and Figure 3. It can be seen that in a laminar regime, the numerical terminal velocity matches the empirical value with a two to three-digit accuracy (Figure 3(a)). The terminal state drag force $F_{D}$ obtained from the Palabos Turret simulation in this regime has less than 0.5% relative difference from the gravitational force. For the higher Reynolds number up to order of 10000, the Palabos Turret simulation results of the terminal state drag force show a relative difference up to 3% from the gravitational force (Figure 3(b)), despite the increase in resolution $N$ from 300 to 400 for the $\mathrm{Re}\geq 1000$ cases. This results in $\sim 137\%$ increase in mesh size, from 81 million to 192 million lattices. To improve the accuracy further, a finer mesh resolution is required to resolve the thinning boundary layer well, but at a much higher computational cost. The terminal velocity from the Palabos Turret simulation $v_{\mathrm{T}}^{D}$ remains below $6\%$ relative difference from its empirical value $v_{\mathrm{T}}^{R}$, which itself has an error up to 6% from the experimental values [10].
The comparison of the $v_{\mathrm{T}}^{D}$ for the $d=0.140$ mm sphere with the terminal velocity obtained from laboratory experiments shows a relative difference lower than 0.2% [69], which further validates the Palabos Turret simulation. In such Palabos Turret simulation, the evolution of the terminal state pressure and shear / viscous force contribution in the total drag force is well resolved and is plotted as a function of the particle Reynolds number in Figure 3(b). It is observed that the pressure contribution to the drag force increases from close to 1/3 in the Stokes regime [65] for $\mathrm{Re}=0.6$ to 97% for $\mathrm{Re}=12000$.
Figure 3: (a) Comparison of terminal velocities $v_{\mathrm{T}}$ of falling spheres at various Reynolds number $\mathrm{Re}$ between the Palabos Turret simulation (Table 3) and an empirical drag model (Eq. 31). For this empirical terminal velocity calculation, the terminal state drag force is considered to be equal to the gravitational force on the sphere. (b) Evolution of the terminal state pressure and shear / viscous force contribution in the total drag force for various $\mathrm{Re}$. The values are the percentage of the pressure force to the total drag force, shown as $p^{*}$ in Table 3.
3.4 Terminal velocities of spheroids and transient oscillations